OpenStudy (anonymous):

Help? Imaginary numbers: (-6)(-7i)(8i)

OpenStudy (mathstudent55):

What is (-6) * (-7) * 8 = ?

OpenStudy (anonymous):

336

OpenStudy (mathstudent55):

Good.

OpenStudy (mathstudent55):

What is i * i = ?

OpenStudy (anonymous):

1?

OpenStudy (mathstudent55):

No. The i of imaginary numbers is defined like this: \(i = \sqrt{-1}\) Since the square of a square root just gives you what is inside the square root, then \(i^2 = (\sqrt{-1})^2 \) or \(i^2 = -1\)

OpenStudy (anonymous):

okay

OpenStudy (mathstudent55):

Now you can think of your problem like this (using the commutative and associative properties as well as knowledge of i): \((-6)(-7i)(8i) \) \(= (-6)(-7)(8) \times (i)(i) \) \(= 336i^2\) \(= 336 \times (-1) \) \(= -336\)

OpenStudy (anonymous):

thanks!

OpenStudy (mathstudent55):

You're welcome.